\begin{thebibliography}{Bou2}

\bibitem[Bor]{Borel}Borel, Armand:
Linear algebraic groups. Second edition. Graduate Texts in Mathematics, 126. Springer-Verlag, New York, 1991.

\bibitem[Bou1]{Bourbaki4}Bourbaki, Nicolas:
\'El\'ements de math\'ematique. Fasc. XXXIV. Groupes et alg\`ebres de Lie. Chapitre IV: Groupes de Coxeter et syst\`emes de Tits. Chapitre V: Groupes engendr\'es par des r\'eflexions. Chapitre VI: syst\`emes de racines.  Actualit\'es Scientifiques et Industrielles, No. 1337 Hermann, Paris, 1968.

\bibitem[Bou2]{Bourbaki7}Bourbaki, Nicolas:
\'El\'ements de math\'ematique. Fasc. XXXVIII: Groupes et alg\`ebres de Lie. Chapitre VII: Sous-alg\`ebres de Cartan, \'el\'ements r\'eguliers. Chapitre VIII: Alg\`ebres de Lie semi-simples d\'eploy\'ees. Actualit\'es Sci. Indust., No. 1364. Hermann, Paris 1975.

\bibitem[CS]{CS}Conway, J.\ H.; Sloane, N.\ J.\ A.:
Sphere packings, lattices and groups. Third edition. With additional contributions by E.\ Bannai, R.\ E.\ Borcherds, J.\ Leech, S.\ P.\ Norton, A.\ M.\ Odlyzko, R.\ A.\ Parker, L.\ Queen and B.\ B.\ Venkov. Grundlehren der mathematischen Wissenschaften, 290. Springer-Verlag, New York, 1999.

%\bibitem[FKS]{FKS}Feger, Robert; Kephart, Thomas W.; Saskowski, Robert J.:
%LieART 2.0 -- A Mathematica Application for Lie Algebras and Representation Theory,
%arXiv:1912.10969.

\bibitem[Gre]{Green}Green, R.\ M. Combinatorics of minuscule representations. Cambridge Tracts in Mathematics, 199. Cambridge University Press, Cambridge, 2013.

\bibitem[Hum]{Hump}Humphreys, James E.:
Introduction to Lie algebras and representation theory. Second printing, revised. Graduate Texts in Mathematics, 9. Springer-Verlag, New York-Berlin, 1978.

\bibitem[Ser]{Serre}Serre, Jean-Pierre:
Groupes alg\'ebriques associ\'es aux modules de Hodge-Tate.  Journes de G\'eom\'etrie Alg\'ebrique de Rennes (Rennes, 1978), Vol. III, pp. 155--188, 
Ast\'erisque, 65, Soc. Math. France, Paris, 1979. 


\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage{amssymb,amscd,amsmath}
\usepackage[all]{xy}
\newtheorem{thm}{Theorem}

\newtheorem{prop}[thm]{Proposition}

\newtheorem{lem}[thm]{Lemma}
\theoremstyle{remark}

\theoremstyle{definition}

\newcommand{\R}{{\mathbb{R}}}
\newcommand{\C}{{\mathbb{C}}}

\newcommand{\Z}{{\mathbb{Z}}}

\newcommand{\End}{\mathrm{End}}

\newcommand{\GL}{\mathrm{GL}}
\newcommand{\SL}{\mathrm{SL}}

\newcommand{\diag}{\mathrm{diag}}

\newcommand{\ad}{\mathrm{ad}}

\renewcommand{\sl}{\mathfrak{sl}}

\newcommand{\g}{\mathfrak g}

\newcommand{\frakt}{\mathfrak t}

\newcommand{\gl}{\mathfrak{gl}}

\newcommand{\Alt}{{\raise 2pt\hbox{$\scriptstyle\bigwedge$}}}

\begin{document}
\title{Disjointness of a simple matrix Lie group and its Lie algebra}

\author{Michael Larsen}
\email{mjlarsen@indiana.edu}
\address{Department of Mathematics\\
    Indiana University \\
    Bloomington, IN 47405\\
    U.S.A.}

    \subjclass{Primary 22E60}

\thanks{The author was partially supported by NSF grant DMS-2001349.}

\begin{abstract}
Let $G$ be a connected closed subgroup of $\GL_n(\C)$ which is simple as a Lie group and which acts irreducibly on $\C^n$.  Regarding both $G$ and its Lie algebra $\g$ as subsets of $M_n(\C)$,
we have $G\cap \g\neq\emptyset$ if and only if $G$ is a classical group and $\C^n$ is a minuscule representation.
\end{abstract}

\maketitle

\begin{center}
\textit{To Pham Huu Tiep on his sixtieth birthday, with admiration}
\end{center}
\vskip 10pt
Let $G$ be a simple complex Lie group with Lie algebra $\g$ and $\rho\colon G\to \GL(V)$ a faithful, irreducible, complex representation of $G$.  Let $\rho_*\colon \g\to \gl(V)$ denote the corresponding Lie algebra representation.
Using the embeddings $G\stackrel{\rho}{\hookrightarrow}\GL(V)\hookrightarrow \End(V)$ and $\g\stackrel{\rho_*}{\hookrightarrow}\gl(V) = \End(V)$, it makes
sense to ask if $G\cap \g$, by which we abbreviate $\rho(G)\cap \rho_*(\g)$, is empty.  As we will see, the answer is usually yes, but the exceptions form a class which has received attention before (see \cite{Serre}).

Recall \cite[VIII, \S7.3]{Bourbaki7} that $V$ is \emph{minuscule} if it is non-trivial and its weights with respect to a choice of maximal torus form a single orbit under the action of the Weyl group.  The minuscule representations are as follows \cite[Appendix]{Serre}:  for type $A_r$, all exterior powers of the natural representation; for $B_r$, the spin representation; for $C_r$, the natural representation; for $D_r$, the natural representation and the semispin representations; for $E_6$, the two $27$-dimensional irreducibles; and for $E_7$, the unique $56$-dimensional irreducible.

The main result of this paper is as follows:

\begin{thm}
\label{main}
Under the above hypotheses, $G\cap \g$ is non-empty if and only if $G$ is of classical type and $V$ is minuscule.
\end{thm}

The rest of the paper proves the theorem.  The main difficulty is to show necessity of the conditions on $(G,V)$.

Let $g\in G$ and $x\in \g$ satisfy $\rho(g) = \rho_*(x)$.
The Jordan decomposition $g = tu$ in $G$ gives the Jordan decomposition $\rho(g) = \rho(t)\rho(u)$ in $\GL(V)$.  In $M_n(\C)$, we can write $\rho(u) = 1+N$, where $N$ is nilpotent and commutes with $\rho(t)$,
so $\rho(t) +\rho(t)N$ is the Jordan-Chevalley decomposition of $\rho_*(x) = \rho(g)$ in $\End(V)$.  By \cite[Chap.~4, Lemma~A]{Hump}, it follows that $\rho(t)\in \rho_*(\g)$.  Without loss of generality, therefore,
we may assume that $g$ is semisimple, and redefining $x$ if necessary,
we still have $\rho(g) = \rho_*(x)$.  Let $T$ be a maximal torus of $G$ containing $g$, and let $\frakt$ denote its Lie algebra.  

Now, $\rho_*(x)$ commutes with $\rho(t)$ for all $t\in T$.  As $\rho_*$ is injective,
$\ad(t)(x) = x$ for all $t\in T$, so $x$ lies in the weight-$0$ space of the adjoint representation of $G$.
In other words, $x \in \frakt$.

As usual, we identify the character group $X^*(T)$ with a subgroup of $\frakt^*$ by means of the diagram
$$\xymatrix{\frakt\ar[r]^{\exp}\ar[d]_{t^*}&T\ar[d]^\chi\\
\C\ar[r]^\exp&\C^\times}$$
relating a character $\chi$ and its corresponding vector $t^*$.
We choose a basis $e_1,\ldots,e_n$ of $V$ consisting of weight vectors for $T$ for characters $\chi_1,\ldots,\chi_n\in X^*(T)$ and let $t^*_1,\ldots,t^*_n$ denote the elements of $\frakt^*$ corresponding to $\chi_1,\ldots,\chi_n$ respectively.
If $a_1,\ldots,a_n\in\Z$ and $\chi_1^{a_1}\cdots \chi_n^{a_n} = 1$, then $a_1t^*_1+\cdots+a_nt^*_n=0$.

\begin{lem}
\label{AP}
If $\chi_i,\chi_j,\chi_k$ are three characters of $V$ such that $\chi_i\chi_k = \chi_j^2$, then $\chi_i(g)=\chi_j(g)=\chi_k(g)$.
\end{lem}

\begin{proof}
As $\chi_i \chi_j^{-2} \chi_k = 1$, we have $t^*_i-2t^*_j+t^*_k=0$.  As $\rho(g) = \rho_*(x)$, we have
$$\chi_i(g) = t^*_i(x),\,\chi_j(g) = t^*_j(x),\,\chi_k(g) = t^*_k(x),$$
so we have
$$4\chi_i(g)\chi_k(g) = 4\chi_j(g)^2 = 4t^*_j(x)^2 = (t^*_i(x)+t^*_k(x))^2 = (\chi_i(g) + \chi_k(g))^2,$$
so $\chi_i(g) = \chi_k(g)$.  Thus,
$$\chi_j(g) = t^*_j(x) = \frac{t^*_i(x)+t^*_k(x)}2 = \frac{\chi_i(g)+\chi_k(g)}2 = \chi_i(g).$$
\end{proof}

\begin{lem}
\label{quadrangle}
If $\chi_i\chi_j = \chi_k\chi_l$ for characters of $V$, then 
\begin{equation}
\label{quadratic}
(z-\chi_i(g))(z-\chi_j(g)) = (z-\chi_k(g))(z-\chi_l(g)).
\end{equation}
\end{lem}

\begin{proof}
The character equality implies
\begin{equation}
\label{equal sum}
t^*_i+t^*_j = t^*_k+t^*_l,
\end{equation}
so
$$\chi_i(g)+\chi_j(g) = t^*_i(x)+t^*_j(x) = t^*_k(x) + t^*_l(x) = \chi_k(g)+\chi_l(g).$$
As $\chi_i(g)\chi_j(g) = \chi_k(g)\chi_l(g)$, equation \eqref{quadratic} follows immediately.
\end{proof}

\begin{lem}
\label{self-dual}
If $V$ is self-dual, then $\chi_i(g)^2+1=0$ for all $\chi_i$.
\end{lem}

\begin{proof}
Let $\chi_i\chi_j=1$.  Then $t^*_i+t^*_j=0$, and
$$\chi_i(g) = t^*_i(x) = -t^*_j(x) = -\chi_j(g) = -\chi_i(g)^{-1}.$$
Thus $\chi_i(g)^2+1=0$.
\end{proof}

\begin{prop}
If $V$ is not a minuscule representation, then $G\cap \g=\emptyset$.
\end{prop}

\begin{proof}
We fix a maximal torus and a Weyl chamber.
Let $\varpi_1,\ldots,\varpi_r$ denote the fundamental characters, and let $\lambda = \sum_i a_i \varpi_i$ denote the dominant weight of $V$.
Let $\alpha$ denote the highest root in the dual root system.  Let $s = \langle \lambda,\alpha^\vee\rangle$.  By \cite[\S21.3]{Hump}, the length of the $\alpha^\vee$-string of weights of $V$  through $\lambda$ is $s+1$.  If $s\ge 2$, then by Lemma~\ref{AP}, $\alpha^\vee(g) = 1$.  By symmetry,
the same is true for all short roots.  Since the root lattice is generated by short roots, it is true for character in the root lattice, so  $\rho(g)$ is an invertible scalar matrix.  However, $\rho_*(x)$ lies in $\sl(V)$ and therefore has trace $0$,
so this is impossible.  Therefore, $s\le 1$.  Now, $\langle \varpi_i,\alpha^\vee\rangle \ge 1$ with equality if and only if $\varpi_i$ is a minuscule weight \cite[Appendix]{Serre}.  Therefore the proposition holds whenever $\lambda$ is a fundamental weight.
On the other hand, if $\sum_i a_i\ge 2$, then
$$\langle \lambda,\alpha^\vee\rangle  = \sum_i a_i \langle \varpi_i,\alpha^\vee\rangle \ge \sum_i a_i \ge 2,$$
and the proposition again holds.
\end{proof}

\begin{prop}
If $G$ is of type $E_6$ and $\dim V = 27$, then $G\cap \g=\emptyset$.
\end{prop}

\begin{proof}
%
If $\{t^*_i(x)\mid 1\le i\le 27\}$ has only $1$ element, then since the set of differences $t_i^*-t_j^*$ span $\frakt^*$, it follows that $x=0$.  This implies $\rho(g)=0$, which is impossible.
If $\{t^*_i(x)\mid 1\le i\le 27\}$ has exactly $2$ elements, then there exists an affine linear function on $\frakt$ which takes the value $0$ and $1$ on every element $t^*_i$ and therefore 
a value in $[0,1]$ on every point of the convex hull of $\{t^*_i\mid 1\le i\le 27\}$ in $X^*(T)\otimes\R\subset \frakt^*$.
This convex hull is the polytope denoted $2_{21}$ in Coxeter's notation, and each of its codimension $1$ faces has $6$ or $10$ vertices \cite[Example 8.5.16]{Green}, so this is not possible.
Therefore, $\{t^*_i(x)\mid 1\le i\le 27\}$ has at least $3$ values.

Endowing $X^*(T)\otimes \R$ with the Weyl-invariant inner product for which the roots have length $\sqrt 2$, we see that the inner product of any weight with any element of the root lattice is integral,
while the inner product of any weight of $V$ with itself is $4/3$ \cite[Planches]{Bourbaki4}.  Therefore, the inner product $\langle t^*_i,t^*_j\rangle$ of any two distinct weights of $V$ is congruent to $1/3$ (mod $1$) and contained in the interval $[-4/3,4/3)$.
The only possible values are therefore $1/3$ and $-2/3$.  In the former case, we say that $t^*_i$ and $t^*_j$ are \emph{skew}.  In the latter case, we say they are \emph{incident}.
(The terminology is motivated by the correspondence
\cite[Example~8.6.4]{Green} between the weights of $V$ and the $27$-lines on a cubic surface.)  For each weight, the number of incident weights is $10$ \cite[Lemma 10.1.6]{Green}.

We claim that Lemma~\ref{quadrangle} implies that $\{t^*_1(x),\ldots,t^*_{27}(x)\}$ has at most $2$ elements.
We prove this by contradiction, assuming that $t^*_i(x)$, $t^*_j(x)$, and $t^*_k(x)$ are pairwise distinct.  There are four possibilities regarding $\{t^*_i,t^*_j,t^*_k\}$ to consider.

\vskip 10pt\noindent
\textbf{Case 1.  No pair of the weights is incident}.  By \cite[Exercise 8.1.11]{Green} and \cite[Lemma 9.2.7]{Green}, the Weyl group of $E_6$ acts transitively on triples of pairwise skew weights.
By \cite[Lemma 10.1.9]{Green}, we may assume without loss of generality that $t^*_i$, $t^*_j$, and $t^*_k$ belong to a $6$-element set of pairwise skew weights and therefore, by \cite[Theorem 10.2.1]{Green}, to one half of a Schl\"afli double six.
In the other half, there are three weights which are linked to each of $t^*_i$, $t^*_j$, and $t^*_k$; choose one of them and denote it $t^*_l$.  
We claim that $-t^*_m-t^*_i$, $-t^*_m-t^*_j$, and $-t^*_m-t^*_k$ are all weights of $V$.  Indeed, they lie in the dual lattice to the root lattice of $E_6$, and each has the same length as a weight of $V$.  This implies that each is either a weight or the negative of a weight  \cite[Chapter 4, (122)]{CS}.  They are actually weights since their inner products with every weight is congruent to $1/3$ (mod $1$) rather than $2/3$.
However,
$$(-t^*_m-t^*_i) + t^*_i = (-t^*_m-t^*_j)+t^*_j = (-t^*_m-t^*_k) + t^*_k.$$
By Lemma~\ref{quadrangle}, this implies that $t^*_i(x)$, $t^*_j(x)$, and $t^*_k(x)$ are all roots of a common quadratic polynomial, which is absurd.

\vskip 10pt\noindent
\textbf{Case 2.  Exactly one pair of the weights is incident}.  We may assume $t^*_i$ and $t^*_j$ are incident.  We have
$$(t^*_i+t^*_j-t^*_k)^2 = \frac 43,$$
By the same reasoning as in Case 1, $t^*_i+t^*_j-t^*_k$ is a weight $t^*_l$,  so \eqref{equal sum} holds.  However, $t^*_k(x)$ is not a root of $(z-t^*_i(x))(z-t^*_j(x))$, contrary to assumption.

\vskip 10pt\noindent
\textbf{Case 3.  Exactly two pairs of the weights are incident}.  Assume that $t^*_i$ and $t^*_j$ are skew, and both are incident to $t^*_k$.
By \cite[Proposition 10.2.7(iii)]{Green},
the number of weights incident to both $t^*_i$ and $t^*_j$ (including $t^*_k$) is $5$, so the number of weights incident to at least one of the two is $15$, including $t^*_k$.
Therefore, among weights not in $\{t^*_i,t^*_j,t^*_k\}$, the number skew to both $t^*_i$ and $t^*_j$ is $10$, and the number incident to $t^*_k$ is $8$.  We conclude that there exists $t^*_m$ which is skew to all three.
Therefore, we may choose two elements of $\{t^*_i,t^*_j,t^*_k\}$ together with $t^*_m$ and obtain three weights, all taking different values in $x$, containing at most a single incident pair.  This is impossible by Cases 1 and 2.

\vskip 10pt\noindent
\textbf{Case 4.  All pairs of the weights are incident}.  The number of incident pairs with one element in $\{t^*_i,t^*_j,t^*_k\}$ and one element in its complement is $24$, so on average each weight in the complement is incident to one element of 
$\{t^*_i,t^*_j,t^*_k\}$.  Let $t^*_m$ be a weight incident to at most one element of $\{t^*_i,t^*_j,t^*_k\}$.  Without loss of generality, we may assume that $t^*_i$, $t^*_j$, and $t^*_m$ all take different values on $x$, and there are at most two incident pairs among them.  This is impossible by Cases 1--3.

\vskip 10pt\noindent
This proves the claim and therefore the proposition.

%
\end{proof}

\begin{prop}
If $G$ is of type $E_7$ and $\dim V = 56$, then $G\cap \g=\emptyset$.
\end{prop}

\begin{proof}
There is a unique irreducible $56$-dimensional representation of $E_7$, and it is therefore self-dual.
By Lemma~\ref{self-dual}, we may assume that $t^*_j(x) = \chi_j(g) \in \{\pm i\}$ for $1\le j\le 56$.
This implies that the set of $t^*_j$ is a union of two hyperplanes, which must therefore bound the convex hull of $\{t^*_j\mid 1\le j\le 56\}$, which is the convex polytope $3_{21}$ in Coxeter's notation.
However, the codimension $1$ faces of $3_{21}$ each have $7$ or $12$ vertices \cite[Exercise 8.5.17]{Green}, so this is impossible.
\end{proof}

This concludes the proof of the only if direction of Theorem~\ref{main}.

\begin{lem}
Let $m\ge 2$ and $j\in [1,m-1]$ be integers, and let $V = \wedge^j \C^m$.  Let $G$ denote the image of $\SL_m(\C)$ in $\GL_n(V)$ via the exterior $j$th power map.  Then $G\cap \g\neq\emptyset$.
\end{lem}

\begin{proof}
Choose $a\in \C$ such that $a^m = \frac j{j-m}$, and let $b = a^{1-m}$.  Let $e_1,\ldots,e_m$ denote the standard basis of $\C^m$.
We fix an ordered basis of $V$ consisting of the vectors of the form $e_{i_1}\wedge\cdots\wedge e_{i_j}$, where $i_1>\cdots>i_j$, taken in lexicographic order.
Thus, the image of $\diag(a,\ldots,a,a^{1-m})\in \SL_m(\C)$ in $\GL(V)$ is
$$\diag(\underbrace{a^j,\ldots,a^j}_{\binom{m-1}j},\underbrace{a^{j-m},\ldots,a^{j-m}}_{\binom{m-1}{j-1}})=\frac{a^j}j\diag(\underbrace{j,\ldots,j}_{\binom{m-1}j},\underbrace{j-m,\ldots,j-m}_{\binom{m-1}{j-1}})\in\g.$$
\end{proof}

\begin{lem}
Let $V$ be a representation of $G$, $T$ a maximal torus of $G$, and $\frakt$ the Lie algebra of $T$.
If $x\in \frakt$ satisfies $t^*_j(x) \in \{\pm i\}$ for all $j$, then $\rho_*(x)\in G\cap \g$.
\end{lem}

\begin{proof}
Choose a basis of $V$ whose $j$th vector belongs to the ($1$-dimensional) $\chi_j$-weight space of $V$.  In terms of this basis, $\rho_*(x)$ is given by the matrix
$\diag(t^*_1(x),\ldots,t^*_n(x))$.  Inside the group $D$ of invertible diagonal matrices, the image $\rho(T)$ is a closed subgroup \cite[Corollary 1.4(a)]{Borel} and it is diagonal, so it is the 
intersection of the kernels of characters of $D$ which vanish on it \cite[Proposition 8.2(c)]{Borel}.  An element of $X^*(D)$ vanishes on $\rho(T)$ if and only if it lies in the kernel of the natural homomorphism
$X^*(D)\to X^*(T)$.  Thus, if $A$ is the set of $n$-tuples $(a_1,\ldots,a_n)\in \Z^n$ such that $\chi_1^{a_1}\cdots\chi_n^{a_n}=1$, then
$$\rho(T) = \{\diag(c_1,\ldots,c_n)\in D\mid c_1^{a_1}\cdots c_n^{a_n}=1\,\forall (a_1,\ldots,a_n)\in A\}.$$

For all $(a_1,\ldots,a_n)\in A$, we have $\sum_j a_jt^*_j = 0$, so $\sum_j a_jt^*_j(x) = 0$.  Writing $t^*_j(x) = \epsilon_j i$, we have $\sum_j a_j \epsilon_j=0$, so
$$\prod_j t^*_j(x)^{a_j} = \prod_j (\epsilon_j i)^{a_j} = \prod_j (i^{\epsilon_j})^{a_j} = i^{\sum_j a_j\epsilon_j} = 1.$$
This implies $\rho_*(x)\in \rho(T)$.

\end{proof}

For minuscule representations when $G$ is of type $B$, $C$, or $D$, such an $x$ always exists; it suffices to find an element of $X^*(T)\otimes \R$ whose inner product with each weight of $V$ is $\pm 1$.
Using the notation of \cite[Planches]{Bourbaki4}, these can be chosen as follows.  For type $B_r$ and highest weight $\varpi_r$, we take $(1,0,\ldots,0)$.  For type $C_r$ and highest weight $\varpi_1$, we take $(1/2,0,\ldots,0)$.
For type $D_r$ and highest weight $\varpi_1$, we take $(1/2,0,\ldots,0)$.  For type $D_r$ and highest weight $\varpi_{r-1}$ or $\varpi_r$, we take $(1,0,\ldots,0)$.

\end{document}